Factorise the following expressions
(l+m)^2 –4lm (Hint: Expand (l+m)^2 first)

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

l+m)2-4lm (Hint: Expand (l+m)2first)
Expand (l+m)2using identity:(x+y)2= x^2+2xy+y^2
(l+m)^2-4lm = l^2+m^2+2lm-4lm
= l^2+m^2-2lm
= (l-m)^2
Using identity: (x- y)^2= x^2 - 2xy+y^2

Thanks